The work of electric field done during the displacement of a negatively charged particle towards a fixed positively charged particle is 9J.  As a result the distance between the charges has been decreased by half.  What work is done by the electric field over the first half of this distance?


Ans: 3J


How we are getting this answer plz explain

Grade:11

1 Answers

arpan chandrashekhar
12 years ago

given:

kQq/(x/2) - kQq/x =9 where x is the intial distance.  . . . 

or kQq/x = 9J

to find is

KQq/(3/4*x) - kQq/x .... as the rdistance b/w the particles becomes 3/4*x as the particle approches by a quater*x

therefore it is kQq/3x =9/3 =3